OpenStudy (trojanpoem):

Problem: y = tan(2arctanx) prove that: (1+x^2)y'' = 2(2y-x)y' I manged to get another solution for it, but I wanna know what's wrong with this solution. or how to complete it. y = 2tan(arctanx)/1-tan^2(arctanx) = 2x/1 - x^2 (1-x^2)y = 2x -2xy + (1-x^2)y' = 2 -2y -2xy' + (-2x)y' + (1-x^2)y'' = 0 (1-x^2)y'' = 2(y - 2x)y'

OpenStudy (michele_laino):

here we have to develop explicitly the indicated computations

OpenStudy (trojanpoem):

Correct, after deriving the given identity , I got to the proof easily. However , The other solution looked easier to me, when I thought of the problem. as I won't have to derive any inverse or trigonometric function.

OpenStudy (trojanpoem):

Do you think that there is a way to turn (1-x^2)y'' = 2(y - 2x)y' into: (1+x^2)y'' = 2(2y-x)y'

OpenStudy (michele_laino):

at first sight I don't know, I think that the best way in order to solve your problem is a direct substitution of the function \(y\) into the ODE

OpenStudy (trojanpoem):

See this too: y = tan^'1 cotx^2 + sin^2cos-1sqrt(x) Using cos^-1 triangle I got that sin^-1 = sqrt(1-x) y = 1/x^2 + 1-x y' = -2x^-3 -1 while I must prove that y' = -2x - 1 I seems like, trying to simplify it, complicates it instead.

OpenStudy (michele_laino):

for example, if we use the chain rule, we can write this: \[\frac{{dy}}{{dx}} = \frac{{dy}}{{dz}} \cdot \frac{{dz}}{{dx}} = \frac{d}{{dz}}\left( {\tan z} \right) \cdot \frac{d}{{dx}}\left( {2\arctan x} \right)\] as you can see, I used this new variable: \(z=2 \arctan x\)

OpenStudy (michele_laino):

and I wrote: \(y=\tan z\)

OpenStudy (michele_laino):

so, after a simple computation, we get: \[\begin{gathered} \frac{{dy}}{{dx}} = \frac{{dy}}{{dz}} \cdot \frac{{dz}}{{dx}} = \frac{d}{{dz}}\left( {\tan z} \right) \cdot \frac{d}{{dx}}\left( {2\arctan x} \right) = \hfill \\ \hfill \\ = \left( {{{\left( {\tan z} \right)}^2} + 1} \right) \cdot \frac{2}{{1 + {x^2}}} = \hfill \\ \hfill \\ = \left( {{{\left( {\tan \left( {2\arctan x} \right)} \right)}^2} + 1} \right) \cdot \frac{2}{{1 + {x^2}}} \hfill \\ \end{gathered} \]

OpenStudy (trojanpoem):

y' = sec^2(2arctan(x)) * 2 * 1/(1+x^2) y' = (1+ tan^2(2arctan(x)) * 2/(1+x^2) (1+x^2)y' = 2(1+y^2) (1+x^2)y'' + 2xy' = 4yy' (1+x^2)y'' = 2(2y - x)y'

OpenStudy (trojanpoem):

The problem is, That confuses me,each time, I have to re-solve the problem twice: 1) with simplification 2) without

OpenStudy (michele_laino):

I don't understand, now we have computed the first derivative, and then we have differentiated the so obtained ODE, so we got the requested ODE, it is a correct procedure

OpenStudy (trojanpoem):

See this y = tan(2arctanx), Why would I differentiate tan, tan inverse when I could differentiate 2x/1+x^2 ?

OpenStudy (michele_laino):

another different procedure is to compute the second derivative, so substituting both the first and the second derivatives into the given ODE, we have to obtain an identity

OpenStudy (michele_laino):

since we can consider that function as a function of function, so we can apply the so called \(chain\;rule\) as I wrote before, we can write this: \(y=\tan z\) and \(z=2 \arctan x\), so if we consider the composed function, we get: \(y= \tan z= \tan ( 2 \arctan x)\) and the chain rule is: \[\frac{{dy}}{{dx}} = \frac{{dy}}{{dz}} \cdot \frac{{dz}}{{dx}}\]

OpenStudy (trojanpoem):

I do chain rule all in one. y' = sec^2(2arctan(x)) * 2 * 1/(1+x^2) y' = (1+ tan^2(2arctan(x)) * 2/(1+x^2) (1+x^2)y' = 2(1+y^2) (1+x^2)y'' + 2xy' = 4yy' (1+x^2)y'' = 2(2y - x)y'

OpenStudy (michele_laino):

correct! your first step is just the \(chain\; rule\)

OpenStudy (trojanpoem):

Why when I simplify it , I don't get what I want -,-

OpenStudy (michele_laino):

I see that all your last steps are correct!

OpenStudy (trojanpoem):

y = 2tan(arctanx)/1-tan^2(arctanx) = 2x/1 - x^2 (1-x^2)y = 2x -2xy + (1-x^2)y' = 2 -2y -2xy' + (-2x)y' + (1-x^2)y'' = 0 (1-x^2)y'' = 2(y - 2x)y'

OpenStudy (michele_laino):

I'm checking your steps, please wait...

OpenStudy (trojanpoem):

Take your time.

OpenStudy (michele_laino):

I'm not able to get the right result, I don't know, I continue to try, and tomorrow I will write my reasonings.

OpenStudy (michele_laino):

oops.. I don't succeed to get the right result...

OpenStudy (trojanpoem):

Thanks so much for your time.

OpenStudy (michele_laino):

:)

OpenStudy (trojanpoem):

If you succeeded , post the result, here and tag me.

OpenStudy (michele_laino):

ok! :)

OpenStudy (michele_laino):

here is my resoning: using the function: \[y = \frac{{2x}}{{1 - {x^2}}}\] I computed both first and second derivatives: \[y' = 2\frac{{1 + {x^2}}}{{{{\left( {1 - {x^2}} \right)}^2}}},\quad y'' = \frac{{4x\left( {3 + {x^2}} \right)}}{{{{\left( {1 - {x^2}} \right)}^3}}}\] using such derivative, we can show that left side = right side of your ODE, and both sides are equal to: \[\frac{{4x\left( {3 + {x^2}} \right)\left( {1 + {x^2}} \right)}}{{{{\left( {1 - {x^2}} \right)}^3}}}\]

OpenStudy (michele_laino):

@TrojanPoem

OpenStudy (trojanpoem):

Nice job, @Michele_Laino .

OpenStudy (michele_laino):

thanks!! :) @TrojanPoem